myininaya (myininaya):

For fun. Prove by induction: \[\frac{d}{dx} |x^n|=nx^{n-2}|x| \text{ for all odd integer } n \ge 1\]

ganeshie8 (ganeshie8):

*

OpenStudy (anonymous):

ermm can i use something else but not induction :P

myininaya (myininaya):

Induction is the preferred method. But I welcome all responses.

OpenStudy (anonymous):

\(\frac{d}{dx} |x^n|=nx^{n-2}|x| \text{ for all odd integer } n \ge 1 \) ok it goes well when n=1 lets this be true for k ( let k be odd) \(\frac{d}{dx} |x^k|=kx^{k-2}|x| \) we need to show for k+2 \(\frac{d}{dx} |x^k+2|= \frac{d}{dx} |x^k x^2|=\frac{d}{dx} |x^k|| x^2|... \text{ since x^2>0 }\\=(kx^{k-2}|x|)x^2+2x(x^k)=2x^{k+1}+kx^k|x|\\ = \)

OpenStudy (anonymous):

hmm something went wrong with my induction xD

OpenStudy (anonymous):

maybe i need to show for 2k+1 not any k

myininaya (myininaya):

Yeah do show for k=1 then assume true it is true for 2k+1 then show it is true for 2k+3 since 2k+3 is the next odd

ganeshie8 (ganeshie8):

hey your proof works just fine except for a silly mistake in the last line

OpenStudy (anonymous):

i thought so .

OpenStudy (anonymous):

|x^k|

myininaya (myininaya):

oh yeah k, k+2 works too where k is odd

OpenStudy (anonymous):

i'll tell u why it works . its induction any order set might work :P

OpenStudy (anonymous):

ordered*

ganeshie8 (ganeshie8):

\(\frac{d}{dx} |x^{k+2}|= \frac{d}{dx} |x^k x^2|=\frac{d}{dx} |x^k|| x^2|... \text{ since x^2>0 }\\=(kx^{k-2}|x|)x^2+2x \color{Red}{|}x^k\color{Red}{|} = kx^k|x| + 2x^k |x| ~~\color{Red}{\star}\\ =(k+2)x^k|x| ~~ \blacksquare \)

OpenStudy (anonymous):

oh yes

ganeshie8 (ganeshie8):

\(\color{Red}{*}\) : when k is odd we have \(|x^k| = x^{k-1}|x|\)

OpenStudy (anonymous):

yes yes i put practices thats why i got confused a bit

OpenStudy (anonymous):

nice question

myininaya (myininaya):

It comes from @idku 's questions for today sorta.

OpenStudy (anonymous):

was gonna say something about induction but na forget it :P

myininaya (myininaya):

What were you going to say earlier before I said induction is the preferred method? Were you just going to do it using the chain rule and all that jazz?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but then realized its hmm so clear in that either hmm

OpenStudy (anonymous):

i should say also not either right ?

myininaya (myininaya):

you are saying 'but then I realized its hmm so clear in that chain rule way' ?

OpenStudy (anonymous):

yes

myininaya (myininaya):

that is what I translated what you said if you are asking about translations

myininaya (myininaya):

but i do not know if either would be the correct word

OpenStudy (anonymous):

haha nvm leave it

OpenStudy (anonymous):

i think also is the right one since induction is also clear :P

myininaya (myininaya):

I think I misunderstood what you said @marki . I'm sorry.

OpenStudy (anonymous):

;)

ganeshie8 (ganeshie8):

I like induction more as i am born biased to induction proofs xD using chain rule : \[\frac{d}{dx}(\left|x^{2k+1}\right|) = \frac{d}{dx}(x^{2k}\cdot \left|x\right|) = 2kx^{2k-1}|x| + x^{2k}\frac{|x|}{x} = (2k+1)x^{2k-1}|x| \]

ganeshie8 (ganeshie8):

this chainrule thing also requires induction to say that it holds for all k i guess

myininaya (myininaya):

We weren't talking about English/grammar at all. You were saying either way works.

OpenStudy (anonymous):

ganesh since u used 2k+1 then no need

ganeshie8 (ganeshie8):

basically all proofs touch induction in someway hmm i could be wrong though with that statement..

OpenStudy (anonymous):

k can be anything the result would odd either way

OpenStudy (anonymous):

now this is what i meant to say about induction >.< it came up after WOP which means to prove for natural in specific this set {1,.....,k,k+1} why we need k and k+1 k for odd\even k+1 foe even\odd

ganeshie8 (ganeshie8):

we can choose to induct on odd integers and define the set with least positive integer accordingly right ?

OpenStudy (anonymous):

so need to be careful , working with induction sometimes is dumb usual they ask u to do it to make u practice more it does not give proper proof , its only for generalization

OpenStudy (anonymous):

yeah u can do anything , but whats the point >.<

ganeshie8 (ganeshie8):

induction works when you have a formula.. it wont help you with deriving the formula though

OpenStudy (anonymous):

yes , means for generalization

ganeshie8 (ganeshie8):

Exactly! generalization(induction) is a valid proof technique. I am not happy about this statement : `it does not give proper proof , its only for generalization`

OpenStudy (anonymous):

hmm ur free to not be happy :P im a fan of wop , i love Archimedean property and glade they made induction of it and happy with D.A

OpenStudy (anonymous):

feeling bored @myininaya any other questions :|

myininaya (myininaya):

not at the moment :( @ganeshie8 is full of them usually

myininaya (myininaya):

I really loved that one question with that one answer that mod 9 question you asked \[5^{2015}+4^{2015} \\ (5-9)^{2015}+4^{2015} \\ (-4)^{2015}+4^{2015} \\ -4^{2015}+4^{2015} \\ 0\]

OpenStudy (anonymous):

haha yeah clever when power is odd

myininaya (myininaya):

that was an answer by @mathmath333 or whatever his name is

OpenStudy (anonymous):

yes i remember

myininaya (myininaya):

i have to go be back for more math fun later

OpenStudy (anonymous):

ok take care bbye !